A varies jointly as B and C,and A=9 when B=5,C=7,find B and A=54,C=10
Answer by Timnewman(323)

Hi dear,
If A varies jointly as B and C,
A=kBC----(1)
Where K is proportionality Constant.
From equ 1,
k=A/BC
When A=9,B=5,C=7
k=9/5*7
k=9/35
Put k=9/35 in equ 1
A=9BC/35(this is the relationship between A,B,C.
when A=54,C=10
Then,
54=(9B*10)/35
54=90B/35
90B=1890
B=21
HOPE THIS HELPS?
Cheers,
Tim.